OpenStudy (anonymous):

What is the remainder when (2x^4 – 3x^3 + x^2 – 3x – 15) ÷ (x – 2) ?

OpenStudy (earthcitizen):

\[2 x^4-3 x^3+x^2-3 x-15 = (3+3 x+x^2+2 x^3)×(-2+x) ,r=-9\]